OpenStudy (lena772):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

ganeshie8 (ganeshie8):

start by finding slope of given line.

OpenStudy (lena772):

Parallel lines have same slope. So I know that it can't be A OR B

OpenStudy (lena772):

y=3x+3 right?

ganeshie8 (ganeshie8):

good observation ! and yes thats right !!

OpenStudy (lena772):

-2=3*4+b -2=12+b -14=b

OpenStudy (lena772):

y=3x-14?

ganeshie8 (ganeshie8):

Correct !